OpenStudy (anonymous):

Find a polynomial f(x) of degree 3 that has the indicated zeros and satisfies the given condition. -4i, 4i, 2; f(4) = -256

OpenStudy (anonymous):

yeh simple

OpenStudy (anonymous):

f(x) =A (x+4i)(x-4i)(x-2)

OpenStudy (anonymous):

okay and if x=4 than it must = -256, right?

OpenStudy (anonymous):

and that's how you find a?

OpenStudy (anonymous):

(x+4i)(x-4i) = x^2 - (-16) = x^2 +16

OpenStudy (anonymous):

f(x) = A (x^2 +16)(x-2)

OpenStudy (anonymous):

when x=4, f= -256

OpenStudy (anonymous):

-256 = 32(2) A

OpenStudy (anonymous):

A= -4 f(x) = -4 ( x^2+16)(x-2)

OpenStudy (anonymous):

Oh I see what I did wrong, I added a negative sign by accident, thanks once again!